OpenStudy (anonymous):

Determine whether the lines intersect. If they do, find the coordinates of the point of intersection. [x, y, z]=[2, 3, 2] + s[3, 4, 10] [x, y, z]=[3, 4, 2] + t[4, 5, 3]

OpenStudy (foolaroundmath):

For line 1 : x = 2+3s, y = 3+4s, z = 2+10s For line 2 : x = 3+4t, y = 4+5t, z = 2+3t Let the two lines intersect. This means that when we solve for s and t, we will get a unique result. 2+3s = 3+4t .... (1) 3+4s = 4+5t .... (2) 2+10s = 2+3t .... (3) From (3), 10s = 3t or s = 3t/10 substituting in (1), we get 9t/10 = 1 + 4t => 31t/10 = -1 => t = -10/31 s = 3t/10 = -3/31 So from (1) and (3), we get s = -3/31, t = -10/31 This value of s and t should also hold true for (2) if there is a point of intersection substituting the values in (2), we get LHS = 3 + 4(-3/31), RHS = 4 + 5(-10/31) LHS = 81/31, RHS = 74/31 As LHS is not equal to RHS, there is no point of intersection